OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the function y = –3x² –12x + 3? (–2, 29) (2, –15) (2, –9) (–2, 15)

OpenStudy (zehanz):

As with the previous one, calculate the x-coordinate with the formula x = -b/(2a). As you can see, a = -3 and b = -12. Put that x-value in y = –3x² –12x + 3 and out comes your y-value of the vertex. It is the same routine every time!

OpenStudy (anonymous):

but I can't do it! It's so hard for me!!!

OpenStudy (zehanz):

Come on, that can't be true. Let's do it together: \(x=-\dfrac{b}{2a}=-\dfrac{-12}{2 \cdot -3}=...\) What number do you get?

OpenStudy (anonymous):

-6

OpenStudy (zehanz):

How much is \(-\dfrac{-12}{-6}\)?

OpenStudy (anonymous):

6

OpenStudy (zehanz):

Try again please...

OpenStudy (zehanz):

it is: \(-\dfrac{12}{6}\)=...

OpenStudy (anonymous):

18???????

OpenStudy (zehanz):

Ignore the first minus-sign. Then you have -12 divided by -6, so that is the same as 12/6. How much is 12/6

OpenStudy (anonymous):

it is 6

OpenStudy (anonymous):

or 2

OpenStudy (zehanz):

6 or 2, only one can be right! So it is 2. Now remember, we still have that first minus-sign, so we've found: x=-2. Now you have to set x=-2 in the equation: y = –3x² –12x + 3 to get y.

OpenStudy (zehanz):

@iloveyooh: YW!

OpenStudy (anonymous):

THANK YOU!